# Domain-Aware Tensor Network Structure Search

### v1

@@ Abstract
Tensor networks (TNs) provide efficient representations of high-dimensional data, yet identification of the optimal TN structures, the so called tensor network structure search (TN-SS) problem, remains a challenge. Current state-of-the-art (SOTA) algorithms are computationally expensive as they require extensive function evaluations, which is prohibitive for real-world applications. In addition, existing methods ignore valuable domain information inherent in real-world tensor data and lack transparency in their identified TN structures. To this end, we propose a novel TN-SS framework, termed the tnLLM, which incorporates domain information about the data and harnesses the reasoning capabilities of large language models (LLMs) to directly predict suitable TN structures. The proposed framework involves a domain-aware prompting pipeline which instructs the LLM to infer suitable TN structures based on the real-world relationships between tensor modes. In this way, our approach is capable of not only iteratively optimizing the objective function, but also generating domain-aware explanations for the identified structures. Experimental results demonstrate that tnLLM achieves comparable TN-SS objective function values with much fewer function evaluations compared to SOTA algorithms. Furthermore, we demonstrate that the LLM-enabled domain information can be used to find good initializations in the search space for sampling-based SOTA methods to accelerate their convergence while preserving theoretical performance guarantees.
@@ 1 Introduction
The exponential increase in the volume and richness of available data has led to the widespread use of multi-way arrays, often represented as higher-order tensors. Tensor network decomposition methods aim to represent higher-order tensors in “super-compressed” formats through smaller-sized components, by effectively capturing cross-modal latent patterns and correlations. These methods have been applied across various fields, including machine learning [1, 2, 3], signal processing [4], computer vision [5, 6], and quantum physics [7, 8]. The success of tensor network decomposition techniques is closely linked to their ability to mitigate the “curse of dimensionality”, which is achieved by decomposing higher-order data into lower-order factors.
However, tensor network (TN) practitioners face significant challenges related to choosing the most appropriate TN structure, the so called tensor network structure search (TN-SS) problem, which has been proven to be NP-hard [9, 10]. TN-SS involves determining the optimal TN parameters, such as TN ranks, TN topology, and TN mode permutations [11].
Existing methods to efficiently resolve the TN-SS problem (see Section 1.1) include Bayesian inference [12], spectrum methods [13], reinforcement learning [14], and discrete [11, 10] and continuous optimization [15], along with greedy search and random sampling. To date, sampling-based methods [16, 11, 10, 17], whose workflows are illustrated in Figure 1a and 1b, have demonstrated the best performance in addressing the TN-SS problem.
However, these methods require large number of evaluations (of the training and testing data), are prone to getting stuck in local minima [10], and lack transparency in their found structure-related parameters. Critically, a large number of evaluations needed to optimize the objective function leads to a high computational cost. We hypothesize that these weaknesses of the current methods are due to the failure to utilize the rich domain information inherent in real-world tensors, such as the mode information. To this end, we ask ourselves:
Can we utilize the intrinsic domain information in tensor data to significantly reduce the number of evaluations required to solve the TN-SS problem, while providing domain-aware explanations for the identified TN solutions?
Figure 1: Comparison of SOTA tensor network structure search (TN-SS) methods. (a) TN-SS algorithms with theoretical guarantees: TNLS [11] & TnALE [10]. (b) tnGPS [17], a prompting pipeline which uses LLMs to generate new TN-SS algorithms. (c) tnLLM (ours), which uses domain information about the tensor data and LLM reasoning to solve the TN-SS problem and generate explanations for the identified TN structure.
@@
[…]
@@
Figure 2: The workflow of the proposed tnLLM framework consists of two key stages commonly adopted by SOTA TN-SS methods [11, 10, 17]: TN-initialization and TN-discovery. The proposed tnLLM efficiently utilises the domain information in tensor data and uses an LLM to guide these two stages.
@@ 1.1 Related work
Compared to traditional tensor decompositions [18, 19, 20, 21, 22], which have pre-defined tensor network structures, the TN-SS problem focuses on finding custom tensor networks, which have been shown to achieve higher parameter efficiency and are an important paradigm of tensor decompositions [10, 11, 16, 17]. Various approaches have been proposed to address the TN-SS problem, including Bayesian inference [12], spectrum methods [13], reinforcement learning [14], program synthesis [23], and continuous optimization [15]. Sampling-based methods [10, 11, 16, 17], which fall under discrete optimization methods, have demonstrated superior performance compared to other approaches to address the TN-SS problem. Among these, TNLS [11] and TnALE [10] achieve state-of-the-art (SOTA) performance, with TnALE accelerating TNLS by slightly reducing the number of evaluations required to optimize the objective function. While both TNLS and TnALE adopt a “local-search” scheme, this acceleration is achieved by the alternating variable update framework proposed in TnALE. However, these methods still face challenges, such as a large numbers of function evaluations required to converge, difficulty in balancing exploration and exploitation, and a lack of transparency in the found TN structures.
More recently, tnGPS [17] has emerged as an approach that uses LLMs to generate sampling-based TN-SS algorithms, demonstrating performance comparable to SOTA methods. Despite being closest to our work, tnGPS focuses on generating TN-SS algorithms based on sampling-based heuristics and does not incorporate any domain information in the tensor data. As such, tnGPS suffers from the same limitations as existing sampling-based methods, including a lack of explainability in the identified structure-related parameters.
@@
The rise of transformer-based LLMs, pretrained on vast text corpora, has demonstrated a remarkable capacity for “reasoning” [24]. This reasoning ability is further enhanced when LLMs are guided by task-specific prompting strategies, such as chain-of thought [25, 26, 27]. These strategies enable LLMs to generate text effectively for tasks such as arithmetic reasoning and factual knowledge retrieval, leading to exceptional performance in complex question-answering and tasks requiring real-world knowledge [28, 29]. These findings suggest that, through pretraining on diverse textual data, LLMs encode rich knowledge about real-world relationships, which they can effectively leverage to perform various downstream tasks [30]. In this work, we utilize the real-world domain knowledge and “reasoning” capabilities of LLMs to directly infer suitable TN structures.
@@ 2 Preliminaries
In this section, we first summarize the notations for TNs and review the TN-SS problem. Then, we provide a brief overview of SOTA sampling-based methods for TN-SS such as TNLS [11], TnALE [10], and tnGPS [17]. Their workflow guides the design of our proposed framework.
@@ 2.1 Tensor network structure search (TN-SS)
We first provide the definition of TN-SS through its application in higher-order data decomposition. An optimal solution to the TN-SS problem aims to find the best trade-off between identifying the most compressed TN structure while preserving the expressivity of the TN [31]. Let $\mathcal{X}\in\mathbb{R}^{I_{1}\times I_{2}\times\cdots\times I_{N}}$ be a non-zero order-$N$ tensor, with its Frobenius norm as $\|\mathcal{X}\|_{F}$. Tensor networks can be represented using the graphical notation [32, 33], whereby a tensor network $G$ is represented using a set of $N$ vertices, $V$, and a set of edges, $E$, i.e., $G=(V,E)$. Each vertex represents a decomposed core tensor, and the closed edges between two core tensors are generalized higher-order matrix multiplications, termed tensor contractions [2]. Closed edges have assigned TN-ranks, $\mathbf{r}\in\mathbb{Z}_{+}^{E}$, which indicate the degree of connections between different pairs of connected vertices. Therefore, the properties of a TN structure can be fully expressed by the combination $(G,\mathbf{r})$.
Similar to SOTA TN-SS methods[16, 10, 11, 17], the discrete optimization problem of TN-SS is formalized as a minimization of the objective function, which is a linear sum of the complexity of the TN structure (e.g., compression rate) and the TN expressivity (e.g., approximation error), and is given by
@@
where $L$ represents the number of tensor samples in the dataset, and the first term $\phi(G,\mathbf{r})$ measures the TN structure complexity. The second term in Equation (1) measures the expressivity of the TN through the relative squared error (RSE) between the original tensors $\{\mathcal{X}_{l}\}_{l=1}^{L}$ and their TN approximations $\left\{TNC\left(\{\mathcal{V}_{l,i}\}_{i=1}^{N};(G,\mathbf{r})\right)\right\}_{l=1}^{L}$, where $TNC(\cdot)$ stands for the tensor contraction operation of the entire tensor network. The pair $(G,\mathbf{r})$ characterizes how the vertices, $\{\mathcal{V}_{l,i}\}_{i=1}^{N}$, are contracted together to approximate the original tensor $\mathcal{X}_{l}$. The coefficient $\lambda$ is a positive non-zero scaling factor which balances the trade-off between the model complexity and model expressivity. Note that [16, 33] pointed out that the TN-SS problem is conveniently equivalent to the TN rank search problem of a fully connected TN [34].
Algorithm 1 : Sampling-based Algorithms for TN-SS
@@
2:$N_{Iter}$triangleright Maximum number of iterations
3:$N_{Sample}$triangleright Number of samples per iteration
4:$P\leftarrow[]$triangleright Historical TN structures
5:$H\leftarrow[]$triangleright Discovered TN structures in each iter.
$\mathcal{F}(\cdot)\leftarrow$ Equation (1) $\triangleright$ Objective function
7:$(G,\mathbf{r})=\textit{Initial TN Structure}$triangleright TN-initialization
@@
14:         $(G,\mathbf{r})\leftarrow(\hat{G},\hat{\mathbf{r}})$.
15:    end if
16:    if Converged then
17:         return $(G,\mathbf{r})$
18:    end if
19:end for
20:Output: $(G,\mathbf{r})$
@@ 2.2 SOTA sampling-based algorithms for TN-SS
Algorithm 1 summarizes the current SOTA algorithms for TN-SS such as [11, 10, 17]. They follow a three-step search process: TN-initialization $\rightarrow$ TN-discovery via sampling in the search neighborhood $\rightarrow$ Updating the center of the search neighborhood. Existing methods ignore the inherent domain information in real-world tensor data, which calls for the development of a framework that can effectively utilize domain knowledge in TN-initialization and TN-discovery to improve performance.
@@
During the TN-initialization stage, the “Task-directive” prompt is used to propose a strong initial TN structure by using the provided domain information. In the TN-discovery stage, the last and best-identified TN structures are used in conjunction with the “Optimization-directive” prompt to guide the LLM in navigating the search space effectively. This iterative process leads to the refinement of the TN structure and yields improved objective function values over successive evaluations.
@@
The interactions between the carefully designed user prompts and the LLM assistant follow an automated structured dialogue [35, 36]. This dialogue-based approach is intuitive for generating stepwise conversational reasoning and particularly useful for iterative tasks such as the TN-SS problem. To further improve the ability of the LLM to perform complex reasoning, we employ the chain-of-thought prompting [25]. This prompting strategy breaks down complex tasks into a series of intermediate reasoning steps and thus enables the model to solve the problem by addressing smaller, simpler sub-tasks sequentially. In the proposed tnLLM framework, three distinctly purposed prompts are employed, each serving a unique role. The full prompts are given in Appendix B.
@@
Figure 4: Structure and components of “Optimization-directive” prompt used for LLM-guided TN-discovery.
@@ 4 Experimental results
In this section, the performance of tnLLM was evaluated against the SOTA sampling-based methods in tensor decomposition tasks in terms of the number of evaluations and objective function value. Our results demonstrate that tnLLM achieves comparable objective function values while delivering significant speed-ups. Moreover, we demonstrate its ability to generate domain-consistent explanations for the identified TN structures. Finally, an ablation study was conducted to evaluate the effectiveness of the proposed prompting pipeline and assess the framework’s robustness across different LLMs.
@@
Given the flexibility of our method to handle tensor data of any order, we evaluated its performance on three types of tensor data with varying sizes and dimensionalities across different domains. In particular, we tested on sets of order-$3$ RGB images data (each tensor sample has size of $144\times 176\times 3$), order-$4$ RGB videos data (each tensor sample has size of $144\times 176\times 3\times 10$), and order-$5$ financial time-series data (each tensor sample has size of $3\times 6\times 3\times 4\times 5$), which, to the best of our knowledge, is the largest dataset in terms of number of samples ever considered in the TN-SS problem. All entries were standardized to values in $[0,1]$, with $80\%$ of each dataset used for training, and the remaining $20\%$ for testing. More details about the data can be found in Appendix C.
@@
In Equation (1), we set $\lambda=10$ and used the same compression ratio function, $\phi$, as in previous sampling-based methods [16], defined as the ratio between the number of parameters in the compressed TN format and the original tensor. The maximum number of evaluations was set to $500$ for the images and videos datasets, and $250$ for the financial time-series dataset. An early stopping criterion with a patience of 5 and a delta of 0 was applied. For all experiments, the LLM model GPT-4o (gpt-4o-2024-08-06) [37] was used, with a temperature setting of 0.2.
@@
In all experiments, we also implemented the three SOTA sampling-based TN-SS algorithms, namely TNLS [11], TnALE [10] and tnGPS [17]. Since the vanilla TNLS is designed to search only for the permutation of a TN, we extended it to fit the settings of TN-SS. For fair comparisons, all baseline methods were evaluated using the same objective function and maximum number of evaluations, with one evaluation defined as a single pass through the entire training/ testing dataset. The full implementation details are provided in Appendix D.
@@
We ran tnLLM five times and examined the mean and standard deviation for both the objective function values and the number of evaluations to demonstrate its robustness. Observe from Table 1 that tnLLM achieved performance on par with SOTA sampling-based algorithms in both training and test objectives across all data types, as measured by the objective function. Importantly, tnLLM minimizes the objective function with significantly fewer evaluations, requiring up to 78×78$\times$78 × fewer evaluations than TNLS, 41×41$\times$41 × fewer than TnALE, and 110×110$\times$110 × fewer than tnGPS.
Remark 4.1.
We used the number of evaluations, rather than runtime, as our primary metric for computational cost comparisons, as it is a standard hardware-invariant metric adopted by SOTA models [11, 10, 17]. Nevertheless, in tnLLM, LLM inference accounted for only 9.41%, 1.19%, and 2.12% of the total runtime on the images, videos, and time-series datasets, respectively. Given the low runtime required by tnLLM, this translates to less than 1 extra evaluation.
Table 1: Performance comparison across different datasets. The values on the left give the lowest training objective function values and their corresponding testing objective function values. The values in [square brackets] give the number of evaluations to first achieve the best training objective function value. For robustness assessment of tnLLM, we report the average and standard deviation of both the objective function value and the number of evaluations across 5 independent runs. For both metrics, a lower value is better. The best values are denoted in bold. The second best values are underlined.
@@
Images Train -0.66 [114] -0.65 [81] -0.66 [438] -0.63±$\pm 0.01$ [4.0$\pm$1.9]
Test -0.47 -0.46 -0.44 -0.48$\pm$0.05
Videos Train -1.64 [484] -1.66 [254] -1.65 [175] -1.63±$\pm 0.01$ [6.2$\pm$3.5]
Test -1.72 -1.72 -1.71 -1.70$\pm$0.02
Time-series Train -0.45 [218] -0.47 [177] -0.39 [38] -0.42±$\pm 0.02$ [5.6$\pm$4.4]
Test -0.43 -0.47 -0.40 -0.41$\pm$0.02
@@
To assess the domain-relevance of the identified TNs, we manually verified the explanations provided for the identified TN structures in all three datasets tested. The explanations of the tnLLM model for the initialized TN structure, with and without domain knowledge, are provided in Figure 6. Observe that without domain information, although the model is inherently capable of reasoning, its lack of domain information about the modes in the real-world tensors causes it to default to random assumptions, likely influenced by its pre-training data, such as assuming a mode as resembling time or frequency. As a result, the explanations it provides are entirely incorrect from a practical perspective and are therefore ineffective in solving TN-SS and guiding TN practitioners.
In contrast, incorporating domain information enables the model to identify the relationships between modes (highlighted in blue in Figure 6) and, more importantly, relate them to real-world principles in vision and finance domains (highlighted in green). The model also shows consistent reasoning in its rank suggestions: when describing a rank as low, medium, or high, it quantitatively selects a value that is coherent relative to other ranks within the same tensor. Moreover, it adjusts these values based on comparisons across different mode pairs, for example by assigning similar ranks to the Height–RGB and Width–RGB modes in the images dataset, and a lower rank to the RGB–Frames mode in the videos dataset. This demonstrates the model’s understanding of the tensor mode relationships based on domain information and supports the validity of the identified TN structure. Consequently, tnLLM offers explanations that are practically useful in helping domain experts to verify and trust the discovered structures. To demonstrate robustness, the complete summarized explanations across three runs for all tested datasets are provided in Appendix E.

### v2

@@ Abstract
Tensor networks (TNs) provide efficient representations of high-dimensional data, yet identification of the optimal TN structures, the so called tensor network structure search (TN-SS) problem, remains a challenge. Current state-of-the-art (SOTA) algorithms solve TN-SS as a purely numerical optimization problem and require extensive function evaluations, which is prohibitive for real-world applications. In addition, existing methods ignore the valuable domain information inherent in real-world tensor data and lack transparency in their identified TN structures. To this end, we propose a novel TN-SS framework, termed the tnLLM, which incorporates domain information about the data and harnesses the reasoning capabilities of large language models (LLMs) to directly predict suitable TN structures. The proposed framework involves a domain-aware prompting pipeline which instructs the LLM to infer suitable TN structures based on the real-world relationships between tensor modes. In this way, our approach is capable of not only iteratively optimizing the objective function, but also generating domain-aware explanations for the identified structures. Experimental results demonstrate that tnLLM achieves comparable TN-SS objective function values with much fewer function evaluations compared to SOTA algorithms. Furthermore, we demonstrate that the LLM-enabled domain information can be used to find good initializations in the search space for sampling-based SOTA methods to accelerate their convergence while preserving theoretical performance guarantees.
@@ 1 Introduction
The exponential increase in the volume and richness of available data has led to the widespread use of multi-way arrays, often represented as higher-order tensors. Tensor network decomposition methods aim to represent higher-order tensors in “super-compressed” formats through smaller-sized components, by effectively capturing cross-modal latent patterns and correlations. These methods have been applied across various fields, including machine learning (Zhe et al., 2015; Cichocki et al., 2016; Malik, 2021), signal processing (Zheng et al., 2023), computer vision (Shakeri & Zhang, 2019; Yamamoto et al., 2022), and quantum physics (Orús, 2019; Felser et al., 2021). The success of tensor network decomposition techniques is closely linked to their ability to mitigate the “curse of dimensionality”, which is achieved by decomposing higher-order data into lower-order factors.
However, tensor network (TN) practitioners face significant challenges related to choosing the most appropriate TN structure, the so called tensor network structure search (TN-SS) problem, which has been proven to be NP-hard (Hillar & Lim, 2013; Li et al., 2023). TN-SS involves determining the optimal TN parameters, such as TN ranks, TN topology, and TN mode permutations (Li et al., 2022).
Existing TN-SS methods solve TN-SS as a purely numerical optimization problem and include approaches such as Bayesian inference (Zeng et al., 2024b), spectrum methods (Chen et al., 2024), reinforcement learning (Meirom et al., 2022), discrete optimization (Li et al., 2022; 2023), and continuous optimization (Zheng et al., 2024). To date, sampling-based methods (Li & Sun, 2020; Li et al., 2022; 2023; Zeng et al., 2024a), whose workflows are illustrated in Figure 1a and 1b, have demonstrated the best performance in addressing the TN-SS problem.
However, these methods require large number of evaluations (of the training and testing data), are prone to getting stuck in local minima, and lack transparency in their found structure-related parameters (Li et al., 2023). Critically, a large number of evaluations needed to optimize the objective function leads to a high computational cost. We hypothesize that these limitations arise from failing to exploit the rich domain information inherent in real-world tensors, such as the mode information. To this end, we ask ourselves:
How can we utilize the intrinsic domain information in tensor data to significantly reduce the number of evaluations required to solve the TN-SS problem, while providing domain-aware explanations for the identified TN solutions?
Figure 1: Comparison of SOTA tensor network structure search (TN-SS) methods. (a) TN-SS algorithms with theoretical guarantees: TNLS (Li et al., 2022) & TnALE (Li et al., 2023). (b) tnGPS (Zeng et al., 2024a), a prompting pipeline which uses LLMs to generate new TN-$S$ algorithms. (c) tnLLM (ours), which uses domain information about the tensor data and LLM reasoning to solve the TN-SS problem and generate explanations for the identified TN structure.
@@
To evaluate the effectiveness of tnLLM, we compare its performance against SOTA TN-SS algorithms on real-world tensor datasets of order-$3,4,\text{and }5$. The experimental results demonstrate that tnLLM delivers performance comparable to current SOTA methods, while requiring significantly fewer function evaluations and providing domain-aware explanations for the identified TN structures. Moreover, we constructed a hybrid algorithm to combine the speed-up benefits of tnLLM with the theoretical guarantees of existing sampling-based approaches. The main contributions of this work are:
@@
Figure 2: The workflow of the proposed tnLLM framework consists of two key stages commonly adopted by SOTA TN-SS methods (Li et al., 2022; 2023; Zeng et al., 2024a): TN-initialization and TN-discovery. The proposed tnLLM efficiently utilises the domain information in tensor data and uses an LLM to guide these two stages.
@@ 1.1 Related work
Compared to traditional tensor decompositions (Hitchcock, 1927; Tucker, 1966; Oseledets, 2011; Zhao et al., 2016; Wu et al., 2022), which have pre-defined tensor network structures, the TN-SS problem focuses on finding custom tensor networks, which have been shown to achieve higher parameter efficiency and are an important paradigm of tensor decompositions (Li et al., 2022; 2023; Li & Sun, 2020; Zeng et al., 2024a). Various approaches have been proposed to address the TN-SS problem, including Bayesian inference (Zeng et al., 2024b), spectrum methods (Chen et al., 2024), reinforcement learning (Meirom et al., 2022), program synthesis (Guo et al., 2025), and continuous optimization (Zheng et al., 2024). Sampling-based methods (Li et al., 2023; 2022; Li & Sun, 2020; Zeng et al., 2024a), which fall under discrete optimization methods, have demonstrated superior performance compared to other approaches in addressing the TN-SS problem. Among these, TNLS (Li et al., 2022) and TnALE (Li et al., 2023) achieve state-of-the-art (SOTA) performance, with TnALE accelerating TNLS by slightly reducing the number of evaluations required to optimize the objective function. While both TNLS and TnALE adopt a “local-search” scheme, this acceleration is achieved by the alternating variable updates proposed in TnALE. However, these methods still face challenges, such as a large number of function evaluations required to converge, difficulty in balancing exploration and exploitation, and a lack of transparency in the found structures.
More recently, tnGPS (Zeng et al., 2024a) has emerged as an approach that uses LLMs to generate sampling-based TN-SS algorithms, demonstrating performance comparable to SOTA methods. Despite being closest to our work, tnGPS focuses on generating TN-SS algorithms based on sampling-based heuristics and does not incorporate any domain information in the tensor data. As such, tnGPS suffers from the same limitations as existing sampling-based methods, including a lack of explainability in the identified structure-related parameters.
@@
The rise of transformer-based LLMs, pretrained on vast text corpora, has demonstrated a remarkable capacity for “reasoning” (Wei et al., 2022a). This reasoning ability is further enhanced when LLMs are guided by task-specific prompting strategies, such as chain-of thought (Wei et al., 2022b; Suzgun et al., 2023; Kojima et al., 2024). These strategies enable LLMs to generate text effectively for tasks such as arithmetic reasoning and factual knowledge retrieval, leading to exceptional performance in complex question-answering and tasks requiring real-world knowledge (Liévin et al., 2023; Singhal et al., 2023). These findings suggest that, through pretraining on diverse textual data, LLMs encode rich knowledge about real-world relationships, which they can effectively leverage to perform various downstream tasks (Choi et al., 2022). In this work, we utilize the real-world domain knowledge and “reasoning” capabilities of LLMs to directly infer suitable TN structures.
@@ 2.1 Tensor network structure search (TN-SS)
We first provide the definition of TN-SS through its application in higher-order data decomposition. An optimal solution to the TN-SS problem aims to find the best trade-off between identifying the most compressed TN structure while preserving the expressivity of the TN (Li & Zhao, 2021). Let $\mathcal{X}\in\mathbb{R}^{I_{1}\times I_{2}\times\cdots\times I_{N}}$ be a non-zero order-$N$ tensor, with its Frobenius norm as $\|\mathcal{X}\|_{F}$. Tensor networks can be represented using the graphical notation (Orús, 2014; Ye & Lim, 2018), whereby a tensor network $G$ is represented using a set of $N$ vertices, $V$, and a set of edges, $E$, i.e., $G=(V,E)$. Each vertex represents a decomposed core tensor, and the closed edges between two core tensors are generalized higher-order matrix multiplications, termed tensor contractions (Cichocki et al., 2016). Closed edges have assigned TN-ranks, $\mathbf{r}\in\mathbb{Z}_{+}^{E}$, which indicate the degree of connectivity between different pairs of connected vertices. Therefore, the properties of a TN structure can be fully expressed by the combination $(G,\mathbf{r})$.
Similar to SOTA TN-SS methods (Li & Sun, 2020; Li et al., 2023; 2022; Zeng et al., 2024a), the discrete optimization problem of TN-SS is formalized as a minimization of the objective function, which is a linear sum of the complexity of the TN structure (e.g., compression rate) and the TN expressivity (e.g., approximation error), and is given by
@@
where $L$ represents the number of tensor samples in the dataset, and the first term $\phi(G,\mathbf{r})$ measures the TN structure complexity. The second term in Equation (1) measures the expressivity of the TN through the relative squared error (RSE) between the original tensors $\{\mathcal{X}_{l}\}_{l=1}^{L}$ and their TN approximations $\left\{TNC\left(\{\mathcal{V}_{l,i}\}_{i=1}^{N};(G,\mathbf{r})\right)\right\}_{l=1}^{L}$, where $TNC(\cdot)$ stands for the tensor contraction operation of the entire tensor network. The pair $(G,\mathbf{r})$ characterizes how the vertices, $\{\mathcal{V}_{l,i}\}_{i=1}^{N}$, are contracted together to approximate the original tensor $\mathcal{X}_{l}$. The coefficient $\lambda$ is a positive non-zero scaling factor which balances the trade-off between model complexity and model expressivity. Note that (Li & Sun, 2020; Ye & Lim, 2018) pointed out that the TN-SS problem is conveniently equivalent to the TN rank search problem of a fully connected TN (Zheng et al., 2021).
Algorithm 1 : Sampling-based Algorithms for TN-SS (Li et al., 2022; 2023; Zeng et al., 2024a)
@@
2: $N_{Iter}$ $\triangleright$ Maximum number of iterations
3: $N_{Sample}$ $\triangleright$ Number of samples per iteration
4: $P\leftarrow[]$ $\triangleright$ Historical TN structures
5: $H\leftarrow[]$ $\triangleright$ Discovered TN structures in each iter.
6: $\mathcal{F}(\cdot)\leftarrow$ Equation (1) $\triangleright$ Objective function
7: $(G,\mathbf{r})=\textit{Initial TN Structure}$ $\triangleright$ TN-initialization
@@
[…]
15:   if Converged then
16:    return $(G,\mathbf{r})$
17: Output: $(G,\mathbf{r})$
@@ 2.2 SOTA sampling-based algorithms for TN-SS
Algorithm 1 summarizes the current SOTA algorithms for TN-SS. They follow a three-step search process: TN-initialization $\rightarrow$ TN-discovery via sampling in the search neighborhood $\rightarrow$ Updating the center of the search neighborhood. Existing methods ignore the inherent domain information in real-world tensor data, which calls for the development of a framework that can effectively utilize domain knowledge in TN-initialization and TN-discovery to improve performance.
@@
During the TN-initialization stage, the “Task-directive” prompt is used to propose a strong initial TN structure by using the provided domain information. In the TN-discovery stage, the last and best-identified TN structures are used in conjunction with the “Optimization-directive” prompt to guide the LLM in navigating the search space effectively. The identified TN structures are fed into an objective function evaluation program to obtain the objective function values. This iterative process leads to the refinement of the TN structure and yields improved objective function values over successive evaluations.
@@
The interactions between the carefully designed user prompts and the LLM assistant follow an automated structured dialogue (OpenAI, 2023; Touvron et al., 2023). This dialogue-based approach is intuitive for generating stepwise conversational reasoning and is particularly useful for iterative tasks such as the TN-SS problem. To further improve the ability of the LLM to perform complex reasoning, we employ the chain-of-thought prompting (Wei et al., 2022b). This prompting strategy breaks down complex tasks into a series of intermediate reasoning steps and thus enables the model to solve the problem by addressing smaller, simpler sub-tasks sequentially. In the proposed tnLLM framework, three distinctly purposed prompts are employed, each serving a unique role. The full prompts are given in Appendix B.
@@
Figure 4: Structure and components of the “Optimization-directive” prompt used for LLM-guided TN-discovery.
@@ 4 Experimental results
In this section, the performance of tnLLM was evaluated against the SOTA sampling-based methods in tensor decomposition tasks in terms of the number of evaluations and objective function value. Our results demonstrate that tnLLM achieves comparable objective function values while delivering significant speed-ups. Moreover, we demonstrate its ability to generate domain-consistent explanations for the identified TN structures. Finally, an ablation study was conducted to evaluate the effectiveness of domain information and assess the framework’s robustness across different LLMs.
@@
Given the flexibility of our method to handle tensor data of any order, we evaluated its performance on three types of tensor data with varying sizes and dimensionalities across different domains. In particular, we tested on datasets of order-$3$ RGB images and order-$4$ RGB videos. We also curated a completely new order-$5$ financial time-series dataset of $142$ tensors to ensure that it was not included in the training data of LLMs. This time-series tensor dataset, to the best of our knowledge, is the largest dataset in terms of number of samples ever considered in the TN-SS problem. All entries were standardized to values in $[0,1]$, with $80\%$ of each dataset used for training, and the remaining $20\%$ for testing. More details about the data can be found in Appendix C.
@@
In Equation (1), we set $\lambda=10$ and used the same compression ratio function, $\phi$, as in previous sampling-based methods (Li & Sun, 2020), defined as the ratio between the number of parameters in the compressed TN format and the original tensor. The maximum number of evaluations was set to $500$ for the images and videos datasets, and $250$ for the financial time-series dataset. An early stopping criterion with a patience of 5 was applied. For all experiments, the LLM model GPT-4o (gpt-4o-2024-08-06) (OpenAI, 2024) was used, with temperature set at 0.2.
@@
In all experiments, we also implemented the three SOTA sampling-based TN-SS algorithms, namely TNLS (Li et al., 2022), TnALE (Li et al., 2023) and tnGPS (Zeng et al., 2024a), and accelerated them with GPUs. Since the vanilla TNLS is designed to search only for the permutation of a TN, we extended it to fit the settings of TN-SS. For fair comparisons, all baseline methods were evaluated using the same objective function and maximum number of evaluations, with one evaluation defined as a single pass through the entire training and testing dataset. The full implementation details are provided in Appendix D.
@@
We ran tnLLM five times and examined the mean and standard deviation for both the objective function values and the number of evaluations to demonstrate its robustness. Observe from Table 1 that tnLLM achieved performance on par with SOTA sampling-based algorithms in both training and test objectives across all data types, as measured by the objective function. Importantly, tnLLM minimizes the objective function with significantly fewer evaluations, requiring up to $78\times$ fewer evaluations than TNLS, $41\times$ fewer than TnALE, and $110\times$ fewer than tnGPS. Consequently, tnLLM achieves runtime reductions of up to 98.3% compared to TNLS, 97.7% compared to TnALE, and 98.1% compared to tnGPS, even after accounting for LLM inference. The full runtime comparisons are provided in Appendix E.
Table 1: Performance comparison across different datasets. The values on the left give the lowest training and corresponding testing objective function values. The values in [square brackets] give the number of evaluations required to first achieve the best training objective function value. For robustness assessment of tnLLM, we report the average and standard deviation of both the objective function value and the number of evaluations across 5 independent runs. For both metrics, a lower value is better. The best values are denoted in bold. The second best values are underlined.
@@
[…]
@@
To assess the domain-relevance of the identified TNs, we manually verified the explanations provided for the identified TN structures in all three datasets tested. The explanations of the tnLLM model for the initialized TN structure, with and without domain knowledge, are provided in Figure 6. Observe that without domain information, although the model is inherently capable of reasoning, its lack of domain information about the modes in the real-world tensors causes it to default to random assumptions, likely influenced by its pre-training data, such as assuming a mode as resembling time or frequency. As a result, the explanations are entirely incorrect from a practical perspective and are therefore ineffective in solving TN-SS.
In contrast, incorporating domain information enables the model to identify relationships between modes (highlighted in blue) and, more importantly, relate them to real-world principles in vision and finance (highlighted in green). The model also shows consistent reasoning in its rank suggestions: when describing a rank as low, medium, or high, it quantitatively selects a value that is coherent relative to other ranks within the same tensor. Moreover, it adjusts these values based on comparisons across different mode pairs, for example by assigning similar ranks to the Height–RGB and Width–RGB modes in the images dataset, and a lower rank to the RGB–Frames mode in the videos dataset. This demonstrates the model’s understanding of tensor mode relationships based on domain information and supports the validity of the identified TN structure. Consequently, tnLLM offers explanations that are practically useful in helping domain experts to verify and trust the discovered structures. Summarized explanations across three runs for all datasets are provided in Appendix F.